(5 pts) A student received an A in English (3 units), a C Education (1 unit). Assume A 4 grade points, B-3 grade points, C-2 grade points, D-1 grade points, and F points. Find the students grade point average , the GPA. , in Psychology (3 units), a B in Biology (4 units), and D in Physical 0 grade
Answer #1

Answer:

units

Grade

Grade points

Total Grade

English

3

A

4

12

Psychology

3

C

2

6

Biology

4

B

3

12

Physical education

1

D

1

1

Total

11

31

GPA = 31/11 = 2.818181818

= 2.82 ( two decimals).